# Non-relativistic transport from frame-indifferent kinetic theory

### v1

@@ VI.2.1 Two-dimensional rotating gas
Let us first consider the case of a two-dimensional gas. In two dimensions there is only one spatial component of the field strength. We define the vorticity scalar as
$B=\frac{1}{2}u^{\rho}\epsilon_{\rho\mu\nu}f^{\mu\nu}\,.$ (107)
The vorticity scalar field, defined in this way, contains both magnetic field and non-inertial effects such as the Coriolis force as seen in the rest frame of the fluid. Note that vorticity can be non-zero even in equilibrium, and therefore $B$ is a zeroth-order quantity. The first-order stress tensor can be written as
@@
The superscript ``$(e)"$ corresponds to a parity-even viscosity tensor, while ``$(o)"$ corresponds to a parity-odd one. Explicitly,
$\eta^{(e)\mu\nu}_{\alpha\beta}=\bar{P}^{\langle\mu}_{\langle\alpha}\bar{P}^{\nu\rangle}_{\beta\rangle},~{}~{}~{}~{}~{}~{}\eta^{(o)\mu\nu}_{\alpha\beta}=u^{\rho}\epsilon_{\rho\lambda\langle\alpha}\bar{P}^{\langle\mu}_{\beta\rangle}h^{\nu\rangle\lambda}\,,$ (110)
$A^{\langle\alpha\beta\rangle}=A^{(\alpha\beta)}-\frac{1}{d}h^{\alpha\beta}g_{\gamma\delta}A^{\gamma\delta}=A^{(\alpha\beta)}-\frac{1}{2}h^{\alpha\beta}g_{\gamma\delta}A^{\gamma\delta}\,.$ (111)
@@
$\lambda^{(e)}=\frac{\tau_{\mathrm{coll}}(2nT)}{1+4\tau_{\mathrm{coll}}^{2}B^{2}}\,,~{}~{}~{}~{}~{}~{}\lambda^{(o)}=\frac{2\tau_{\mathrm{coll}}^{2}B(2nT)}{1+4\tau_{\mathrm{coll}}^{2}B^{2}}\,.$ (112)
Regarding the heat current $q^{\mu}$,
$q^{\mu}=-\kappa^{(e)}\partial^{\mu}T-\kappa^{(o)}u^{\rho}\epsilon_{\rho\lambda\nu}h^{\mu\lambda}\partial^{\nu}T$ (113)
@@
$\kappa^{(e)}=\frac{\tau_{\mathrm{coll}}}{1+\tau_{\mathrm{coll}}^{2}B^{2}}\frac{2nT}{m}\,,~{}~{}~{}~{}~{}~{}~{}~{}\kappa^{(o)}=\frac{\tau_{\mathrm{coll}}^{2}B}{1+\tau_{\mathrm{coll}}^{2}B^{2}}\frac{2nT}{m}\,.$ (114)
@@ VI.2.2 Three-dimensional rotating gas
Let us now consider the more complicated case of a 3D gas. In order to express the viscosity covariantly, we first introduce the vorticity vectors with lower index ($B_{\mu}$) and upper index ($B^{\mu}$), their norm $|B|$, and the corresponding unit vectors $b^{\mu}$ and $b_{\mu}$ as follows:
@@
Note that the vectors above are purely spatial, i.e. $\tau_{\mu}b^{\mu}=u^{\mu}b_{\mu}=0$. They can be used to form two spatial projection operators, $R^{\mu}_{\nu}$ longitudinal to $b^{\mu}$ and $Q^{\mu}_{\nu}$ transverse to $b^{\mu}$:
@@
The viscosity tensor $\eta^{\mu\nu}_{\alpha\beta}$ in the absence of bulk viscosity can be written as a sum of five terms [73, 2]
@@
$\begin{split}\eta^{(1e)\mu\nu}_{\alpha\beta}&=R^{\langle\mu}_{\langle\alpha}R^{\nu\rangle}_{\beta\rangle}\,,~{}~{}~{}~{}~{}~{}\eta^{(2e)\mu\nu}_{\alpha\beta}=R^{\langle\mu}_{\langle\alpha}Q^{\nu\rangle}_{\beta\rangle}\,,~{}~{}~{}~{}~{}~{}\eta^{(3e)\mu\nu}_{\alpha\beta}=Q^{\langle\mu}_{\langle\alpha}Q^{\nu\rangle}_{\beta\rangle}\,,\\ \eta^{(1o)\mu\nu}_{\alpha\beta}&=u^{\rho}b^{\sigma}\epsilon_{\rho\sigma\lambda\langle\alpha}R^{\langle\mu}_{\beta\rangle}h^{\nu\rangle\lambda}\,,~{}~{}~{}~{}~{}~{}\eta^{(2o)\mu\nu}_{\alpha\beta}=u^{\rho}b^{\sigma}\epsilon_{\rho\sigma\lambda\langle\alpha}Q^{\langle\mu}_{\beta\rangle}h^{\nu\rangle\lambda}\,.\end{split}$ (118)
Let us work in cylindrical coordinates $(z,\rho,\phi)$ such the $z$ direction is aligned with the vorticity vector, i.e., $B_{\mu}dx^{\mu}=Bdz$. In the case of a rotating gas, $B=2\Omega$ and $z$ is the direction of the axis of rotation. Due to magnetic field anisotropy, viscosity coefficients are different for the different directions:
$\lambda^{(1e)}$ $=\tau_{\mathrm{coll}}(2nT)\,,~{}~{}~{}~{}~{}~{}\lambda^{(2e)}=\frac{\tau_{\mathrm{coll}}(2nT)}{1+\tau_{\mathrm{coll}}^{2}B^{2}}\,,~{}~{}~{}~{}~{}~{}\lambda^{(3e)}=\frac{\tau_{\mathrm{coll}}(2nT)}{1+4\tau_{\mathrm{coll}}^{2}B^{2}}\,,$ (119)
$\lambda^{(1o)}$ $=\frac{\tau_{\mathrm{coll}}^{2}B(2nT)}{1+\tau_{\mathrm{coll}}^{2}B^{2}}\,,~{}~{}~{}~{}~{}~{}\lambda^{(2o)}=\frac{2\tau_{\mathrm{coll}}^{2}B(2nT)}{1+4\tau_{\mathrm{coll}}^{2}B^{2}}\,.$ (120)
In Appendix A we also calculate the heat current $q^{\mu}$:
@@
$\kappa^{(1e)}=\tau_{\mathrm{coll}}\left(\frac{5nT}{2m}\right)\,,~{}~{}~{}~{}~{}~{}~{}~{}\kappa^{(2e)}=\frac{\tau_{\mathrm{coll}}}{1+\tau_{\mathrm{coll}}^{2}B^{2}}\left(\frac{5nT}{2m}\right)\,,~{}~{}~{}~{}~{}~{}~{}~{}\kappa^{(o)}=\frac{\tau_{\mathrm{coll}}^{2}B}{1+\tau_{\mathrm{coll}}^{2}B^{2}}\left(\frac{5nT}{2m}\right)\,.$ (122)
@@ VII Frame-dependence paradox and its resolution
In the scientific literature a great deal of controversy about the principle of frame invariance can be found. Let us present the matter of the debate in the language of the NC kinetic theory. Assume for concreteness a two-dimensional gas with zero shear, $\sigma^{\mu\nu}=0$, but non-zero gradient of temperature: $\partial_{\mu}T\neq 0$. The traditional approach used, e.g., in Ref. [6] is to perform the gradient expansion in the frame comoving with the observer, rather than in the frame comoving with the fluid. In other words, the gradient expansion is performed using $v^{\mu}$ of the observer and the corresponding $h_{\mu\nu}$ and $F_{\mu\nu}$ instead of the Milne-invariant $u^{\mu}$, $g_{\mu\nu}$ and $f_{\mu\nu}$. At the first order, this is equivalent to using the Milne-boost-dependent vorticity scalar $B=\frac{1}{2}v^{\rho}\epsilon_{\rho\mu\nu}F^{\mu\nu}$ instead of the Milne-invariant one in Eq. (107). Let us see what effect it has on the constitutive relations. Firstly, for an observer with zero vorticity, $B=0$, the heat current is given by
@@
Now let us change the frame of reference to one rotating with frequency $\Omega$, as in Section V. This produces vorticity $B=2\Omega$, and following Eq. (113) we then have
@@
The problem is that $q_{1}^{\mu}\neq q_{2}^{\mu}$, despite the seemingly covariant notation for $q^{\mu}$. This apparent contradiction led the author of Ref. [6] to conclude that constitutive relations depend on the observer’s frame of reference. Taking this point of view, however, has paradoxical consequences: the equation of motion for temperature, which in the absence of shear reads
@@
will change form depending on the observer reference frame, even though $T$ is a frame-invariant scalar quantity.
@@
The source of the problem, then, is the arbitrariness in the choice of $v^{\mu}$. The paradox can be resolved by fixing $v^{\mu}$ in terms of some distinguished timelike (in the sense $\tau_{\mu}v^{\mu}=1$) vector field. For generic gases, there exists only one such field, namely the fluid velocity $u^{\mu}$. Therefore, one can obtain a truly covariant kinetic theory by setting $v^{\mu}=u^{\mu}$, which is the approach pursued in the present paper. This is also consistent with the solution proposed in Refs. [15, 16, 17], which suggest that constitutive relations should be calculated in the reference frame corotating with the fluid.
We would now like to focus on two attempts to address the problem of frame invariance in Newtonian physics by employing a covariant language. Firstly, Matolcsi in [20] proposed a mathematical formalism to describe a non-relativistic four-dimensional spacetime and formulated the appropriate Boltzmann equation. This formalism was then used by Matolcsi and Gruber in [24] to perform the gradient expansion in a truly frame-invariant way. The issue with this approach, however, is that the geometry of Matolcsi’s spacetime does not enjoy the Milne boost invariance property. This is equivalent to the spacetime possessing a distinguished vector $v^{\mu}$, which is tied to the motion of “inertial observers”. This allows the authors of [24] to perform the gradient expansion in $P^{\nu}_{\mu}\nabla_{\nu}$, which they call the “absolute spatial derivative”. In a NC geometry, however, the object $P^{\nu}_{\mu}\nabla_{\nu}$ is not Milne-invariant, and there is no reason to distinguish certain observers as “inertial”; the only distinguished observer is the one comoving with the fluid, i.e., one for which $v^{\mu}=u^{\mu}$. Additionally, the construction of [24] does not admit a curved spacetime and there is no mention of the gauge field. We note that in the modern formulation of geometry such spacetimes, without the boost symmetry, are referred to as “Aristotelian”.
@@
Secondly, Frewer in [35] considered the issue of frame invariance in the context of curved manifolds that are NC geometries in our formalism. His solution consists of expressing constitutive equations using tensors called in our notation $u^{\mu}$ and $g_{\mu\nu}$. Despite these tensors being Milne-invariant and in this sense the essence of our solution, Frewer makes no mention of Milne boosts. In consequence, the reason behind the failure of the previous formulations remains a mystery. In addition, unlike the present paper, Ref. [35] does not in fact develop kinetic theory, and consequently a covariant implementation of the gradient expansion procedure is also lacking.
@@ VIII Discussion
The findings of this study significantly advance our understanding of the kinetic theory in non-relativistic gravitational settings, facilitated by the NC framework. The introduction of Milne-invariant data is pivotal in establishing a robust theoretical foundation for analyzing gases and fluids in gravitational fields. As a consequence we are able to resolve a long-standing puzzle in kinetic theory that concerns the covariant formulation. In the Newton-Cartan framework it is clearly visible that the Milne boost symmetry is spontaneously broken by the presence of matter, which fixes the derivative expansion unambiguously. This observation justifies our prescription to formulate frame-indifferent kinetic theory. Without the explicit Milne boost symmetry it is impossible to understand the resolution of the paradox. Previous attempts of this problem have missed this symmetry and thus have not been able to provide a clear physical explanation of the paradox.
@@
Our work not only advances the theoretical understanding of kinetic theory in curved spacetime but also opens up new avenues for experimental investigation in ultracold atomic systems. The integration of geometric potentials and quantum vortices provides a richer framework for exploring the dynamics of superfluids, and the experimental feasibility of studying these systems in microgravity environments further enhances the potential impact of our findings. Recent experimental advances in Bose-Einstein condensates on the International Space Station now allow the creation of ultracold atomic bubbles [77, 78]. This provides a promising possibility to investigate a bubble-trapped superfluid experimentally. Motivated by this experimental progress, there has been a renewed interest in the dynamics of few-body vortices on curved surfaces [79, 80, 81, 82], offering different perspectives on the mathematical treatment of point vortex dynamics on curved surfaces.
@@
The covariant formulation of kinetic theory allows one to systematically study transport in non-inertial frames and under the influence of gravity. Helioseismology has unveiled the internal rotation profile of the Sun [83]. The inner radiative zone exhibits rigid rotation, whereas the outer convective zone displays a complex differential rotation profile. Separating these two zones is a thin layer known as the solar tachocline, which marks the transition from rigid to differential rotation. Current observations indicate that the tachocline’s thickness and position remain relatively stable over time, and it is characterized by strong toroidal magnetic fields. The dynamics of the solar tachocline are a focal point of extensive research [84]. Typically, the plasma dynamics within this thin layer are approximated by the shallow water magnetohydrodynamic equations [85]. In this approximation the curvature effects are neglected. Our frame-indifferent kinetic theory approach enables the systematic incorporation of geometric effects into the magnetohydrodynamic equations and shedding further light into the dynamics of Sun.

### v2

@@ VI.2.1 Two-dimensional rotating gas
Let us first consider the hydrodynamic flow of a two-dimensional gas around an equilibrium state with non-vanishing vorticity. Therefore, we define the vorticity pseudoscalar
$B=\frac{1}{2c}u^{\rho}\epsilon_{\rho\mu\nu}f^{\mu\nu}\,.$ (109)
The vorticity scalar field, defined in this way, contains a contribution of the Coriolis force as measured in the rest frame of the fluid. In that case, we can identify $B=2\Omega$ with $\Omega$ the angular frequency, according to Eq. (91). The first-order stress tensor can be written as
@@
The superscript (e)``(e)" corresponds to a parity-even viscosity tensor, while (o)``(o)" corresponds to a parity-odd one. Explicitly,
$\eta^{(e)\mu\nu}_{\alpha\beta}=\bar{P}^{\langle\mu}_{\alpha}\bar{P}^{\nu\rangle}_{\beta},~~~~~~\eta^{(o)\mu\nu}_{\alpha\beta}=c^{-1}u^{\rho}\epsilon_{\rho\lambda\alpha}\bar{P}^{\langle\mu}_{\beta}h^{\nu\rangle\lambda}\,.$ (112)
@@
$\lambda^{(e)}=\frac{2\mathfrak{p}\,\tau_{\mathrm{coll}}}{1+(2\tau_{\mathrm{coll}}B)^{2}}\,,~~~~~~\lambda^{(o)}=\frac{4\mathfrak{p}\,B\,\tau_{\mathrm{coll}}^{2}}{1+(2\tau_{\mathrm{coll}}B)^{2}}\,.$ (113)
Regarding the ”heat” current $q^{\mu}$,
$q^{\mu}=-\kappa^{(e)}\partial^{\mu}T-\kappa^{(o)}u^{\rho}\epsilon_{\rho\lambda\nu}h^{\mu\lambda}\partial^{\nu}T,$ (114)
@@
$\kappa^{(e)}=\frac{\epsilon+\mathfrak{p}}{\rho}\frac{n\,\tau_{\mathrm{coll}}}{1+(\tau_{\mathrm{coll}}B)^{2}}\,,~~~~~~~~\kappa^{(o)}=\frac{\epsilon+\mathfrak{p}}{\rho}\frac{\tau_{\mathrm{coll}}^{2}B}{1+(\tau_{\mathrm{coll}}B)^{2}}\,.$ (115)
The constitutive relations in a frame with observer velocity $v^{\mu}$ up to the first order in the derivative expansion read
$\mathcal{P}_{\mu}$ $=\rho\,u_{\mu}\,,\qquad T_{\mu\nu}=\mathfrak{p}\,h_{\mu\nu}-\eta_{\mu\nu\alpha\beta}\sigma^{\alpha\beta}\,$, (116)
$\tilde{\mathcal{E}}^{\mu}$ $=\left(\epsilon+\frac{\rho}{2}u^{2}\right)u^{\mu}+\mathfrak{p}\,P^{\mu}_{\nu}\,u^{\nu}-\kappa^{(e)}\partial^{\mu}T-\kappa^{(o)}u^{\rho}\epsilon_{\rho\lambda\nu}h^{\mu\lambda}\partial^{\nu}T-\eta^{\mu\nu}_{\alpha\beta}u_{\nu}\sigma^{\alpha\beta}\,.$ (117)
@@ VI.2.2 Three-dimensional rotating gas
Let us now consider the more complicated case of a 3D gas. In order to express the constitutive relations covariantly, we first introduce the vorticity vectors with lower index ($B_{\mu}$) and upper index ($B^{\mu}$), their norm $|B|$, and the corresponding unit vectors $b_{\mu}$ and $b^{\mu}$, as follows:
@@
Note that the vectors above are purely spatial, i.e., $\tau_{\mu}b^{\mu}=u^{\mu}b_{\mu}=0$. They can be used to form two spatial projection operators, $R^{\mu}_{\nu}$ longitudinal to $b^{\mu}$ and $Q^{\mu}_{\nu}$ transverse to $b^{\mu}$:
@@
The viscosity tensor $\eta^{\mu\nu}_{\alpha\beta}$ in the absence of bulk viscosity can be written as a sum of five terms, [74, 2]
@@
$\begin{split}\eta^{(1e)\mu\nu}_{\alpha\beta}&=R^{\langle\mu}_{\alpha}R^{\nu\rangle}_{\beta}\,,~~~~~~\eta^{(2e)\mu\nu}_{\alpha\beta}=R^{\langle\mu}_{\alpha}Q^{\nu\rangle}_{\beta}\,,~~~~~~\eta^{(3e)\mu\nu}_{\alpha\beta}=Q^{\langle\mu}_{\alpha}Q^{\nu\rangle}_{\beta}\,,\\ \eta^{(1o)\mu\nu}_{\alpha\beta}&=u^{\rho}b^{\sigma}\epsilon_{\rho\sigma\lambda\alpha}R^{\langle\mu}_{\beta}h^{\nu\rangle\lambda}\,,~~~~~~\eta^{(2o)\mu\nu}_{\alpha\beta}=u^{\rho}b^{\sigma}\epsilon_{\rho\sigma\lambda\alpha}Q^{\langle\mu}_{\beta}h^{\nu\rangle\lambda}\,.\end{split}$ (121)
Let us work in cylindrical coordinates $(z,\rho,\phi)$ where the $z$ direction is aligned with the vorticity vector, i.e., $B_{\mu}dx^{\mu}=Bdz$. In the case of a rotating gas, $B=2\Omega$ and $z$ is the direction of the axis of rotation. Due to magnetic field anisotropy, viscosity coefficients are different for the different directions:
$\lambda^{(1e)}$ $=2\mathfrak{p}\,\tau_{\mathrm{coll}}\,,~~~~~~\lambda^{(2e)}=\frac{2\mathfrak{p}\,\tau_{\mathrm{coll}}}{1+\tau_{\mathrm{coll}}^{2}B^{2}}\,,~~~~~~\lambda^{(3e)}=\frac{2\mathfrak{p}\,\tau_{\mathrm{coll}}}{1+4\tau_{\mathrm{coll}}^{2}B^{2}}\,$, (122)
$\lambda^{(1o)}$ $=\frac{2\mathfrak{p}\,B\,\tau_{\mathrm{coll}}^{2}}{1+\tau_{\mathrm{coll}}^{2}B^{2}}\,,~~~~~~\lambda^{(2o)}=\frac{4\mathfrak{p}\,B\,\tau_{\mathrm{coll}}^{2}}{1+4\tau_{\mathrm{coll}}^{2}B^{2}}\,.$ (123)
The heat current $q^{\mu}$ reads
@@
$\kappa^{(1e)}=\frac{\epsilon+\mathfrak{p}}{\rho}n\tau_{\mathrm{coll}}\,,~~~~~~~~\kappa^{(2e)}=\frac{\epsilon+\mathfrak{p}}{\rho}\frac{n\tau_{\mathrm{coll}}}{1+\tau_{\mathrm{coll}}^{2}B^{2}}\,,~~~~~~~~\kappa^{(o)}=\frac{\epsilon+\mathfrak{p}}{\rho}\frac{nB\tau_{\mathrm{coll}}^{2}}{1+\tau_{n\mathrm{coll}}^{2}B^{2}}\,.$ (125)
@@ VII Frame-dependence paradox and its resolution
In the scientific literature, a great deal of controversy about the principle of frame invariance can be found. Let us present the matter of the debate in the language of the NC kinetic theory. Assume for concreteness a two-dimensional gas with zero shear, $\sigma^{\mu\nu}=0$, but non-zero gradient of temperature: $\partial_{\mu}T\neq 0$. The traditional approach used, e.g., in Ref. [6], is to perform the gradient expansion in the frame comoving with the observer, rather than in the frame comoving with the fluid. In other words, the gradient expansion is performed using $v^{\mu}$ of the observer, together with the corresponding $h_{\mu\nu}$ and $F_{\mu\nu}$, instead of the fluid-frame quantities $u^{\mu}$, $g_{\mu\nu}$ and $f_{\mu\nu}$. At the first order, this is equivalent to using the Milne-boost-dependent vorticity scalar $B=\frac{1}{2}v^{\rho}\epsilon_{\rho\mu\nu}F^{\mu\nu}$ instead of the Milne-invariant one defined in Eq. (109). Let us see what effect it has on the constitutive relations. Firstly, for an observer with zero vorticity, $B=0$, the heat current is given by
@@
Now let us change the frame of reference to one rotating with frequency $\Omega$, as in Section V. This gives a non-zero vorticity $B=2\Omega$, and, following Eq. (114), we then have
@@
The problem is that $q_{1}^{\mu}\neq q_{2}^{\mu}$, despite the seemingly covariant notation for $q^{\mu}$. This apparent contradiction led the author of Ref. [6] to conclude that constitutive relations depend on the observer’s frame of reference. Taking this point of view, however, leads to a paradoxical situation: the heat transport equation in the absence of shear,
@@
appears to change its form under different choices of observer frame, even though $T$ is a frame-invariant scalar.
@@
The source of the problem, then, is that the usual Newtonian approach to non-relativistic physics implicitly introduces the observer velocity $v^{\mu}$, which is in principle arbitrary, into the gradient expansion procedure. The way in which $v^{\mu}$ enters the formulas is difficult to spot in a Newtonian formulation, and only becomes evident in a Newton-Cartan treatment. The paradox can be resolved by replacing $v^{\mu}$ with an unambiguously defined timelike (in the sense $\tau_{\mu}v^{\mu}=1$) vector field. Since for generic gases it is only possible to define one such field, namely the fluid velocity $u^{\mu}$, it becomes clear that a truly covariant kinetic theory can be obtained after replacing $v^{\mu}$ with $u^{\mu}$ in the constitutive relations, which is equivalent to setting the observer velocity $v^{\mu}=u^{\mu}$ in the conventional Newtonian approach. This is also consistent with the solution proposed in Refs. [15, 16, 17], which suggest that constitutive relations should be calculated in the reference frame corotating with the fluid.
Our argument rests on the following chain of reasoning:
  • Coordinate-independence is a foundational principle in both relativistic and non-relativistic physics. In the latter case, the Newton–Cartan framework includes not only diffeomorphism invariance but also Milne boost symmetry as a gauge redundancy.
  • In the presence of a thermal fluid, the fluid velocity field $u^{\mu}$ selects a preferred element in the Milne equivalence class, spontaneously breaking this symmetry. This breaking is not an artefact of coordinate choice, but rather introduces fixed background structures—such as $u^{\mu}$ and $g_{\mu\nu}$—that do not transform under Milne boosts and are physically meaningful. These backgrounds play an essential role in defining thermodynamic and hydrodynamic quantities.
  • The constitutive relations must be built from Milne-invariant data, even if Milne boost symmetry is spontaneously broken. This is analogous to constructing observables in gauge theories from gauge-invariant combinations of fields. The comoving frame naturally arises as a consequence of this symmetry-breaking pattern, not merely as a pragmatic or “obvious” choice.
  • Thus, our claim is that the Milne gauge structure and its spontaneous breaking by the thermal state are essential ingredients for a conceptually complete and consistent theory of non-relativistic, covariant kinetic theory. This insight appears to have been largely overlooked in previous treatments, which either worked in fixed frames or implicitly broke Milne invariance without recognizing it as a gauge symmetry.
We now turn our attention to two distinct approaches aimed at addressing the issue of frame invariance in Newtonian physics through the use of covariant formulations.
Firstly, Matolcsi in [20] introduced a mathematical framework describing the flat non-relativistic four-dimensional spacetime. This formalism was then used by Matolcsi and Gruber in [24], where they developed kinetic theory that they argued to be frame-invariant. Matolcsi’s approach restores frame indifference by eliminating any reference to specific observers or coordinate systems. The underlying affine structure of this spacetime provides a canonical flat, torsion-free covariant derivative defined by directional derivatives, so that parallel transport is path-independent and compatibility conditions uniquely determine the clock form and spatial metric. In this sense, Matolcsi’s construction can be regarded as a particular realization of a Newton–Cartan spacetime with closed clock form and flat connection. While this formulation successfully resolves frame indifference in the flat case, it does not directly address the origin of apparent violations in coordinate-based implementations, nor does it offer a route to incorporating curvature or more general geometric structures. The framework developed here aims to close this gap by extending frame-indifferent formulations to settings where coordinate descriptions are indispensable and where nontrivial Newton–Cartan geometries may arise, thereby enabling applications in both flat and curved non-relativistic spacetimes.
@@
Secondly, Frewer in [35] considered the issue of frame invariance in the context of curved manifolds that are NC geometries in our formalism. His solution consists of expressing constitutive equations using tensors called in our notation $u^{\mu}$ and $g_{\mu\nu}$. While expressing the constitutive equations using these tensors is fundamental to our solution of the paradox as well, Frewer makes no mention of the observer velocity $v^{\mu}$ or the Milne symmetry. In consequence, the reason behind the failure of the previous formulations remains a mystery. In fact, unlike the present paper, Ref. [35] is not concerned with developing a covariant kinetic theory, and consequently a covariant implementation of the gradient expansion procedure is also lacking.
@@ VIII Discussion
The findings of this study significantly advance our understanding of the kinetic theory in non-relativistic gravitational settings, facilitated by the NC framework. In both relativistic and non-relativistic settings, physical theories must remain observer-independent, even when spacetime symmetries are broken. In Newton–Cartan geometry, fluids in equilibrium select a preferred frame, spontaneously breaking Milne boost invariance. This breaking introduces fixed background fields, such as the fluid velocity, which do not transform under the broken symmetry. Though the equations remain covariant, these backgrounds encode physically meaningful symmetry breaking, making its effects observable. A consistent use of observer-independent data is pivotal in establishing a robust theoretical foundation for analyzing gases and fluids in gravitational fields. As a consequence we are also able to resolve a long-standing puzzle in kinetic theory that concerns the covariant formulation. Without explicitly enforcing the Milne boost symmetry, and pointing out where the traditional approach to kinetic theory fails to observe it, it is impossible to understand the resolution of the frame-indifference controversy. Previous attempts to solve this problem have missed this symmetry and thus have not been able to provide a clear physical explanation of the paradox.
@@
Our work not only advances the theoretical understanding of kinetic theory in curved spacetime, but also opens up new avenues for experimental investigation in ultracold atomic systems. The integration of geometric potentials and quantum vortices provides a rich framework for exploring the dynamics of superfluids, and the experimental feasibility of studying these systems in microgravity environments further enhances the potential impact of our findings. Recent experimental advances in Bose-Einstein condensates on the International Space Station allow the creation of ultracold atomic bubbles [78, 79]. This provides a promising possibility to investigate a bubble-trapped superfluid experimentally. Motivated by this experimental progress, there has been a renewed interest in the dynamics of few-body vortices on curved surfaces [80, 81, 82, 83], offering different perspectives on the mathematical treatment of point vortex dynamics on curved surfaces.
@@
The covariant formulation of kinetic theory allows one to systematically study transport in non-inertial frames and under the influence of gravity. Helioseismology has unveiled the internal rotation profile of the Sun [84]. The inner radiative zone exhibits rigid rotation, whereas the outer convective zone displays a complex differential rotation profile. Separating these two zones is a thin layer known as the solar tachocline, which marks the transition from rigid to differential rotation. Current observations indicate that the tachocline’s thickness and position remain relatively stable over time, and it is characterized by strong toroidal magnetic fields. The dynamics of the solar tachocline are a focal point of extensive research [85]. Typically, the plasma dynamics within this thin layer are approximated by the shallow water magnetohydrodynamic equations [86]. In this approximation, the curvature effects are neglected. Our frame-indifferent kinetic theory approach enables a systematic incorporation of geometric effects into the magnetohydrodynamic equations and shedding further light on the dynamics of Sun.